# Efficient Dual-domain Image Dehazing with Haze Prior Perception

### v1

@@ Abstract
Transformer-based models exhibit strong global modeling capabilities in single-image dehazing, but their high computational cost limits real-time applicability. Existing methods predominantly rely on spatial-domain features to capture long-range dependencies, which are computationally expensive and often inadequate under complex haze conditions. While some approaches introduce frequency-domain cues, the weak coupling between spatial and frequency branches limits the overall performance. To overcome these limitations, we propose the Dark Channel Guided Frequency-aware Dehazing Network (DGFDNet), a novel dual-domain framework that performs physically guided degradation alignment across spatial and frequency domains. At its core, the DGFDBlock comprises two key modules: 1) the Haze-Aware Frequency Modulator (HAFM), which generates a pixel-level haze confidence map from dark channel priors to adaptively enhance haze-relevant frequency components, thereby achieving global degradation-aware spectral modulation; 2) the Multi-level Gating Aggregation Module (MGAM), which fuses multi-scale features through diverse convolutional kernels and hybrid gating mechanisms to recover fine structural details. Additionally, a Prior Correction Guidance Branch (PCGB) incorporates a closed-loop feedback mechanism, enabling iterative refinement of the prior by intermediate dehazed features and significantly improving haze localization accuracy, especially in challenging outdoor scenes. Extensive experiments on four benchmark haze datasets demonstrate that DGFDNet achieves state-of-the-art performance with superior robustness and real-time efficiency. Code is available at: https://github.com/Dilizlr/DGFDNet.
Index Terms:
Image Dehazing, Frequency Learning, Dark-channel Guidance.
@@ I Introduction
Images captured in hazy conditions often suffer from reduced contrast, weakened color saturation, and loss of high-frequency details. These degradations negatively affect downstream tasks such as object detection [1, 2] and semantic segmentation [3, 4]. Single-image dehazing aims to recover a clear scene from a hazy input. However, it remains a fundamentally ill-posed inverse problem due to the unknown and spatially varying distribution of haze.
Early methods [5, 6, 7, 8] relied on handcrafted priors to estimate transmission maps and restore scene radiance. These methods are effective in simple scenes but often fail in complex real-world scenarios. With the rise of deep learning, CNN-based methods [9, 10, 11, 12] have significantly improved reconstruction quality via end-to-end learning. However, their limited receptive fields restrict the modeling of global dependencies, which are crucial for restoring structural coherence and missing details. Recent Transformer-based architectures [13, 14, 15] have demonstrated strong capabilities in capturing long-range dependencies via self-attention mechanisms, further advancing dehazing performance. Yet, their quadratic complexity severely restricts their applicability in real-time scenarios.
To improve the efficiency of global context modeling, recent studies [16, 17, 18, 19] have explored incorporating frequency-domain information into the dehazing process. This is motivated by the observation that haze mainly distorts the real and magnitude components of the Fourier spectrum, while the phase and imaginary parts largely preserve structural integrity (Fig. 2(a)). Moreover, local frequency alterations can induce global spatial effects (Fig. 2(b)), making frequency-based processing inherently more efficient for capturing long-range dependencies. However, existing spatial-frequency fusion methods typically employ loosely coupled dual-branch designs without explicit degradation alignment across domains. Such weak coupling limits effective information exchange and reduces robustness in complex haze scenarios.
@@
To overcome these limitations, we propose the Dark Channel Guided Frequency-aware Dehazing Network (DGFDNet), a novel dual-domain framework that explicitly aligns haze degradation cues across spatial and frequency domains. By integrating haze localization in the spatial domain with targeted modulation in the frequency domain, DGFDNet effectively captures global context while maintaining computational efficiency, achieving a favorable balance between performance and practicality (see Fig. 1).
Figure 2: (a) Investigation of the degradation characteristics of hazy images in the frequency domain by separately swapping the phase and imaginary components of hazy and clean images. (b) Demonstration of the global impact of frequency domain operations in the spatial domain by modifying local regions of the amplitude in hazy images.
DGFDNet is composed of a cascade of multi-scale DGFDBlocks, each containing two key modules: the Haze-Aware Frequency Modulator (HAFM) and the Multi-level Gating Aggregation Module (MGAM). HAFM generates a pixel-level haze confidence map using dark channel prior features for precise haze localization. Guided by this spatial prior, it adaptively modulates frequency-domain components to emphasize haze-relevant spectra while suppressing irrelevant information, thereby improving global degradation modeling and long-range context capture. MGAM focuses on local detail restoration by employing multi-scale convolutional kernels to capture diverse spatial patterns. Its core structure is a hybrid gating mechanism with hierarchical interaction that adaptively aggregates multi-scale features, enhancing reconstruction fidelity and fine-grained texture recovery.
Although the dark channel prior offers a coarse haze estimate, it often fails in complex outdoor scenes. To remedy this, we introduce the Prior Correction Guidance Branch (PCGB), which employs a closed-loop feedback mechanism to iteratively refine the prior using intermediate dehazed features. This progressive correction improves haze localization accuracy and significantly enhances the robustness and adaptability of the network under challenging real-world conditions.
By combining these components, each DGFDBlock jointly refines spatial and frequency-domain features, enhancing the robustness and effectiveness of DGFDNet. Our main contributions are summarized as follows:
1) We propose DGFDNet, a novel dual-domain dehazing framework that explicitly aligns degradation cues across spatial and frequency domains. Each DGFDBlock integrates HAFM for haze-aware frequency band modulation to model global context and MGAM for fine-grained detail restoration via adaptive multi-scale fusion.
2) We design PCGB, which employs a closed-loop feedback mechanism to iteratively refine the guidance of the dark channel prior based on intermediate dehazed features. This dynamic correction significantly improves haze localization and model robustness in complex outdoor scenes.
3) Extensive experiments on both real-world and synthetic datasets demonstrate that DGFDNet achieves state-of-the-art dehazing performance while maintaining competitive computational efficiency.
@@ II-A Prior-based dehazing methods
Early dehazing approaches relied on handcrafted priors to constrain the solution space, typically derived from statistical analysis and empirical observations of natural images. Representative priors include the haze-line prior [20], color-line prior [21], dark channel prior (DCP)[5], color attenuation prior[6], and color ellipsoidal prior [8]. Among these, DCP is the most widely adopted. It assumes that in haze-free images, at least one color channel has near-zero intensity in most local patches, whereas haze increases this intensity, thereby enabling haze distribution estimation.
However, these priors often struggle in complex real-world scenes due to mismatches between their assumptions and actual conditions. For instance, DCP tends to fail in sky regions, where naturally low intensity contradicts its assumptions, resulting in notable transmission estimation errors.
II-B CNNs-based dehazing methods
@@
CNN-based methods leverage deep convolutions to learn features from large-scale data, offering greater adaptability than handcrafted priors. Early models such as DehazeNet [9], AOD-Net [22], and MSCNN [23] laid the foundation for CNN-based dehazing. Subsequent works integrated attention mechanisms to enhance feature representation, including GridDehazeNet [24] with multi-scale fusion and FFA-Net [25] with pixel and channel attention. HFMCA-Net [26] further improves multi-scale feature learning through hierarchical fusion and mixed convolution attention, while more recent approaches such as TFFD-Net [27] and DEA-Net [28] focus on fine detail recovery through task-specific modules.
Despite their success, CNNs still struggle to model long-range dependencies due to limited receptive fields. Deepening the network can help, but it increases computational cost and latency, limiting real-time deployment.
II-C Transformer-based dehazing methods
@@
Transformers [29, 30, 31, 32, 33, 34] have attracted significant attention in image restoration tasks for their strong global modeling capabilities, including dehazing. Dehamer[35] is among the first to refine local CNN features using Transformer-based global representations. DehazeFormer[14] adapts the Swin Transformer with task-specific designs, while STH[36] adopts a multi-branch architecture to handle varying haze densities. DehazeDCT[37] integrates deformable convolutions to better address spatially variant haze, and MB-TaylorFormer V2 [38] introduces a Taylor series-inspired approximation to reduce Transformer complexity.
While Transformer-based models excel at capturing long-range dependencies, they suffer from two major limitations: quadratic complexity that grows with image resolution, and difficulty in restoring fine-grained local details essential for high-quality reconstruction.
@@
Figure 3: Overview of DGFDNet. (a) It includes a Dehazing Main Branch and Prior Correction Guidance Branch (PCGB), both with a three-scale symmetric design and inter-stage information exchange. The core module, DGFDBlock, The core module, DGFDBlock, incorporates (b) HAFM for global context modeling and (c) MGAM for local information modeling.
II-D Frequency-domain dehazing methods
Recent studies have explored frequency-domain techniques for single-image dehazing. Nehete et al.[16] propose a two-stage network that processes amplitude and phase separately. WSAMF-Net[39] integrates spatial and frequency features using a wavelet-based spatial attention module. Cui et al.[40] introduce a dual-domain selection mechanism to extract key features from both domains. However, they only suppress low-frequency components without global spectral modulation, and their spatial attention fails to accurately localize degraded regions. Yu et al.[17] propose a dual-guided framework to fuse spatial and frequency cues, but their amplitude-driven phase reconstruction may propagate errors when amplitude features are corrupted.
These methods demonstrate the promise of frequency-aware restoration. However, most of them rely on shallow fusion strategies with loosely coupled dual-branch designs. They lack explicit alignment of degradation information across domains, which limits the integration of complementary features.
@@ III Method
As shown in Fig. 3 (a), DGFDNet consists of a dehazing main branch and a prior correction guidance branch. The main branch adopts a U-shaped encoder-decoder with three symmetric stages, each containing multiple DGFDBlocks. The number of blocks and channels varies across stages to handle different resolutions. Strided convolution enables downsampling, while upsampling combines $1\times 1$ convolution and pixel shuffle. Skip connections facilitate feature transfer between encoder and decoder stages. The guidance branch extracts haze-related features from the dark channel prior to assist HAFM in haze perception and receives feedback from MGAM for dynamic refinement. It follows the same downsampling and upsampling structure as the main branch to ensure consistency.
Given a hazy image $I\in\mathbb{R}^{3\times H\times W}$, a $3\times 3$ convolution extracts shallow features $X_{f}\in\mathbb{R}^{C\times H\times W}$, while the dark channel prior is computed using [5]. The shallow features and dark channel prior are fed into the main and guidance branches, respectively, which interact bidirectionally to generate restored features. A final $3\times 3$ convolution maps these features back to the image space, producing a residual image $I_{r}\in\mathbb{R}^{3\times H\times W}$. The dehazed output is then obtained as $I_{c}=I+I_{r}$.
@@ III-A1 Motivation
The spectral characteristics of haze degradation (Fig. 2) suggest that frequency-domain processing is more effective than spatial-domain methods in separating haze from background content. However, its inherent global modulation reduces sensitivity to subtle degradations, which can distort fine details. To solve this, many methods [16, 17] adopt dual-branch architectures that process spatial and frequency domains separately before fusing information via interaction modules. But the loose coupling often limits effective information exchange and fails to fully leverage the complementary strengths of both domains.
In contrast, the proposed HAFM alternates processing between spatial and frequency domains to maximize their synergy while maintaining efficiency. It first focuses on haze-affected regions in the spatial domain to extract structural priors that guide frequency-domain restoration. Specifically, we use the dark channel guidance to generate a spatial attention map that more accurately localizes degraded areas and quantifies their severity than traditional spatial attention. Crucially, this attention mechanism inherently modulates haze-related frequency bands, enabling frequency-domain processing to rapidly capture haze features and perform precise modulation. Such explicit alignment of degradation cues strengthens global modeling and preserves fine details for high-quality restoration.
@@ III-A2 Pipeline
Given the input feature $X_{f}^{i,j-1}$ and the dark channel-guided feature $X_{d}^{i,j-1}$ for the $j$-th DGFDBlock at the $i$-th stage, the HAFM processing can be formulated as:
@@
where $\mathrm{BN}$ represents Batch Norm, $\hat{X}_{f}^{i,j}$ is the residual output feature of HAFM, and $\hat{X}_{d}^{i,j}$ is the intermediate dark channel-guided feature.
Spatial Modulation. As illustrated in Fig. 3(b), HAFM first performs spatial modulation. The dark channel-guided feature $X_{d}^{i,j-1}$ is passed through two consecutive $1\times 1$ convolutions with a GELU activation in between, producing the intermediate dark channel feature $\hat{X}_{d}^{i,j}$. Both convolutions serve to compress channels. Then, $\hat{X}_{d}^{i,j}$ is processed with a Sigmoid function to produce the spatial attention map $M_{sa}$. This process is expressed as:
@@
Meanwhile, the normalized input feature $\mathrm{BN}(X_{f}^{i,j-1})$ is processed by a $1\times 1$ convolution to produce $X_{m}^{i,j}$, for inter-channel information exchange. To recover high-frequency details lost in hazy regions, a $3\times 3$ depth-wise convolution is applied to $X_{m}^{i,j}$. The result is then element-wise multiplied by the dark channel-guided spatial attention map $M_{sa}$ and added to $X_{m}^{i,j}$, yielding the final spatially modulated feature:
@@
Frequency Modulation. Global frequency modulation is then applied to the spatially modulated feature $X_{spatial}^{i,j}$. First, the feature undergoes a Fast Fourier Transform (FFT, $\mathcal{F}$) to obtain its real and imaginary components, $X_{\mathcal{R}}^{i,j}$ and $X_{\mathcal{I}}^{i,j}$. Each part is processed independently using two $1\times 1$ convolutions with a GELU activation in between, functioning like a frequency-domain MLP. This design preserves spectral structure while avoiding cross-component interference. Then, the modulated parts are transformed back to the spatial domain via the inverse FFT (IFFT, $\mathcal{F}^{-1}$), yielding the frequency-modulated feature $X_{frequency}^{i,j}$. The entire process is formulated as:
@@
Finally, the spatially and frequency-modulated features are concatenated along the channel dimension and subsequently fused using a $1\times 1$ convolution followed by a GELU activation to generate the final output:

### v2

@@ Abstract
Transformers offer strong global modeling for single-image dehazing but come with high computational costs. Most methods rely on spatial features to capture long-range dependencies, making them less effective under complex haze conditions. Although some integrate frequency-domain cues, weak coupling between spatial and frequency branches limits their performance. To address these issues, we propose the Dark Channel Guided Frequency-aware Dehazing Network (DGFDNet), a dual-domain framework that explicitly aligns degradation across spatial and frequency domains. At its core, the DGFDBlock consists of two key modules: 1) Haze-Aware Frequency Modulator (HAFM), which uses dark channel priors to generate a haze confidence map for adaptive frequency modulation, achieving global degradation-aware spectral filtering. 2) Multi-level Gating Aggregation Module (MGAM), which fuses multi-scale features via multi-scale convolutions and a hybrid gating mechanism to recover fine-grained structures. Additionally, the Prior Correction Guidance Branch (PCGB) incorporates feedback for iterative refinement of the prior, improving haze localization accuracy, particularly in outdoor scenes. Extensive experiments on four benchmark datasets demonstrate that DGFDNet achieves state-of-the-art performance with improved robustness and real-time efficiency. Code is available at: https://github.com/Dilizlr/DGFDNet.
@@ I Introduction
Hazy images suffer from contrast degradation, color distortion, and detail loss, posing challenges for downstream tasks such as object detection [1, 2] and semantic segmentation [3, 4]. Single-image dehazing seeks to recover a clear scene from a hazy input, but it is an ill-posed problem due to the spatially varying and unknown haze distribution.
Early methods [5, 6, 7, 8] employ handcrafted priors to estimate haze distribution, effective in simple scenes but underperforming in more complex ones. CNN-based methods [9, 10, 11, 12] improve reconstruction quality through end-to-end learning, but their limited receptive fields hindered global dependency modeling, which is essential for restoring structural consistency and lost details. Transformer-based models [13, 14, 15] excel at capturing long-range dependencies via self-attention mechanisms, advancing dehazing performance. However, their quadratic complexity restricts real-time applications.
Recent work [16, 17, 18, 19] has explored integrating frequency-domain information to improve global context modeling efficiency. This is driven by two key insights: 1) Haze primarily distorts the real and magnitude components of the Fourier spectrum, while the phase and imaginary parts retain structural integrity (Fig. 2 (a)), making frequency learning ideal for capturing haze properties. 2) Local frequency changes induce global spatial effects (Fig. 2 (b)), suggesting that frequency-based processing is more efficient for modeling long-range dependencies. However, existing spatial-frequency fusion methods typically use loosely coupled dual-branch designs, lacking explicit degradation alignment, which limits information exchange and robustness in complex haze conditions.
@@
To overcome these limitations, we propose the Dark Channel Guided Frequency-aware Dehazing Network (DGFDNet), a novel dual-domain framework that explicitly aligns haze degradation cues across spatial and frequency domains, enhancing robustness and achieving a compact design. By combining spatial haze localization with targeted frequency-domain modulation, DGFDNet effectively captures global context while maintaining computational efficiency, striking a balance between performance and practicality (Fig. 1).
Figure 2: (a) Investigation of haze degradation in the frequency domain by separately swapping phase and imaginary components between hazy and clean images.(b) Demonstration of the global impact of frequency-domain operations on spatial domain by modifying local amplitude regions in hazy images.
DGFDNet consists of multi-scale DGFDBlocks, each containing two key modules: the Haze-Aware Frequency Modulator (HAFM) and the Multi-level Gating Aggregation Module (MGAM). Guided by the physically grounded dark channel prior, HAFM generates a pixel-level haze confidence map and selectively modulates frequency components related to haze, enabling global degradation modeling. Building on this, MGAM refines spatial representations through multi-scale convolutions and a hybrid gating mechanism, enhancing fine-grained perception and recovering high-frequency details.
While the dark channel prior provides a coarse haze estimate, it struggles in complex outdoor scenes. To tackle this, we introduce the Prior Correction Guided Branch (PCGB), which not only delivers haze degradation cues to HAFM but also receives refinement feedback from MGAM, iteratively correcting the prior information. This closed-loop feedback mechanism enforces consistent alignment of multi-stage, multi-domain degradation information, improving haze localization accuracy and robustness in challenging real-world conditions.
Each DGFDBlock jointly optimizes spatial and frequency-domain features under PCGB guidance, enabling efficient global context modeling while preserving local details. Our main contributions are summarized as follows:
1) We propose DGFDNet, a dual-domain dehazing framework that explicitly aligns spatial and frequency degradation cues. Each DGFDBlock integrates HAFM for haze-aware spectral modulation and MGAM for detail enhancement, balancing global context modeling and local restoration.
2) We design PCGB, which employs a closed-loop feedback mechanism to iteratively refine the dark channel guidance. This dynamic correction strategy significantly improves haze localization and model robustness in complex outdoor scenes.
3) Extensive experiments on real-world and synthetic datasets show that DGFDNet achieves state-of-the-art dehazing performance with competitive computational efficiency.
@@ II Related Work
Prior-based dehazing methods. Early approaches employ handcrafted priors to constrain the solution space, including haze-lines [20], color ellipsoidal priors [8], color-lines [21], dark channel (DCP) [5], and color attenuation [6]. Among these, DCP is the most widely used, assuming that in haze-free images, at least one color channel is nearly zero in most local regions, while haze increases intensity, enabling haze distribution estimation. However, these priors struggle in complex scenes due to discrepancies with real-world conditions.
@@
CNNs-based dehazing methods. CNNs automatically learn features from large datasets, offering better adaptability than priors. Early models such as DehazeNet [9], AOD-Net [22], and MSCNN [23] laid the foundation for deep learning in dehazing. Later approaches, like GridDehazeNet [24] and FFA-Net [25], integrate attention mechanisms to enhance feature extraction. MSAFF-Net [26] further enhances multi-scale learning with spatial attention and feature fusion, while more recent approaches like DFR-Net [27] and DEA-Net [28] focus on fine detail recovery through task-specific modules. However, CNNs struggle with long-range dependencies, and increasing depth introduces higher computational costs.
@@
Transformer-based dehazing methods. Due to strong global modeling capabilities, Transformers [29, 30, 31, 32, 33] are popular in image restoration, including dehazing. Dehamer [34] first integrates Transformer, refining local CNN features with global representations. DehazeFormer [14] adapts the Swin Transformer for dehazing, while STH [35] employs a multi-branch structure to process haze information at different intensities. DehazeDCT [36] incorporates deformable convolutions to handle non-uniform haze. MB-TaylorFormer V2 [37] introduces a Taylor series-inspired approximation to reduce Transformer complexity. Despite these advances, Transformer-based methods still suffer from quadratic complexity as resolution increases and struggle to restore local details.
@@
Figure 3: Overview of DGFDNet. (a) It includes a Dehazing Main Branch and Prior Correction Guidance Branch (PCGB), both with a three-scale symmetric design and inter-stage information exchange. The core module, DGFDBlock, integrates (b) HAFM for global context modeling and (c) MGAM for local information modeling.
Frequency-domain dehazing methods. Recent studies have explored frequency-domain techniques for dehazing. Nehete et al. [16] propose a two-stage network for amplitude and phase processing. Cui et al. [38] introduce a dual-domain selection mechanism but focus mainly on suppressing low frequencies without full global spectral modulation. Their spatial attention also struggles with precise haze localization. Yu et al. [17] introduce a dual-guided framework, but their phase reconstruction may propagate errors if amplitude features are corrupted. These methods demonstrate the potential of frequency-aware dehazing, but often rely on loosely coupled designs that lack explicit degradation alignment across domains, which limits the integration of complementary features.
@@ III Method
As shown in Fig. 3 (a), DGFDNet consists of a dehazing main branch and a prior correction guidance branch. The main branch uses a U-shaped encoder-decoder with three symmetric stages, each containing multiple DGFDBlocks. Downsampling and upsampling use standard techniques, with skip connections transferring features between the encoder and decoder. The guidance branch leverages the dark channel prior to assist HAFM in haze perception, while receiving feedback from MGAM for dynamic prior refinement. It follows the same downsampling and upsampling structure as the main branch to ensure consistency.
Given a hazy image $I\in\mathbb{R}^{3\times H\times W}$, a $3\times 3$ convolution extracts shallow features $X_{f}\in\mathbb{R}^{C\times H\times W}$, while the dark channel prior is computed as in [5]. The shallow features and the prior are fed into the main and guidance branches, respectively, which interact bidirectionally to generate restored features. A final $3\times 3$ convolution produces a residual image $I_{r}\in\mathbb{R}^{3\times H\times W}$, and the dehazed output is $I_{c}=I+I_{r}$.
@@ III-A1 Motivation
The spectral characteristics of haze degradation (Fig. 2) suggest that frequency-domain processing is more effective than spatial-domain methods for separating haze from the background. However, global modulation in the frequency domain may reduce sensitivity to subtle degradations, leading to distortion of fine details. To address this, many methods [16, 17] adopt dual-branch architectures that process spatial and frequency domains separately before merging them. However, such loosely coupled designs often limit the exchange of complementary information.
In contrast, our HAFM alternates between spatial and frequency-domain processing to maximize synergy. The spatial domain first identifies haze-affected regions, providing clear guidance for frequency-domain restoration. Specifically, the dark channel prior generates a spatial attention map that localizes degraded areas and quantifies their severity. This attention mechanism inherently enhances haze-related frequency bands, allowing frequency-domain processing to efficiently capture haze features and apply precise modulation. This explicit alignment of degradation cues enhances global context modeling while enabling fine-grained degradation perception for comprehensive dehazing.
@@ III-A2 Pipeline
For the $j$-th DGFDBlock at the $i$-th stage, given the input feature $X_{f}^{i,j-1}$ and the dark channel-guided feature $X_{d}^{i,j-1}$, HAFM processes them as:
@@
where $\mathrm{BN}$ represents BatchNorm, $\hat{X}_{f}^{i,j}$ is the residual output, and $\hat{X}_{d}^{i,j}$ is the intermediate dark channel-guided feature.
Spatial Modulation. As shown in Fig. 3 (b), the dark channel-guided feature $X_{d}^{i,j-1}$ is passed through two $1\times 1$ convolutions with GELU activation, producing the intermediate feature $\hat{X}_{d}^{i,j}$, which is then processed by a Sigmoid function to produce the spatial attention map $M_{sa}$:
@@
Meanwhile, the input feature $\mathrm{BN}(X_{f}^{i,j-1})$ undergoes a $1\times 1$ convolution to generate $X_{m}^{i,j}$ for inter-channel interaction. A $3\times 3$ depth-wise convolution is used to enhance high-frequency areas. The result is element-wise multiplied by $M_{sa}$ and added to $X_{m}^{i,j}$ to produce the spatially modulated feature:
@@
Frequency Modulation. The spatially modulated feature $X_{spatial}^{i,j}$ is then subjected to frequency modulation. First, the feature undergoes Fast Fourier Transform (FFT, $\mathcal{F}$) to extract its real and imaginary components, $X_{\mathcal{R}}^{i,j}$ and $X_{\mathcal{I}}^{i,j}$, which are processed separately using two $1\times 1$ convolutions with GELU activation. The modulated components are subsequently converted back to the spatial domain via inverse FFT (IFFT, $\mathcal{F}^{-1}$), resulting in the frequency-modulated feature $X_{frequency}^{i,j}$:
@@
Finally, the spatially and frequency-modulated features are concatenated and fused using a $1\times 1$ convolution and GELU activation to produce the final output: